The owner of a newsstand in a college community estimates the weekly demand for a certain magazine as follows.

Quantity Demanded 10 11 12 13 14 15

Probability 0.20 0.15 0.30 0.10 0.05 0.20

Find the number of issues of the magazine that the newsstand owner can expect to sell per week. (Enter your answer to two decimal places.) issues
Answer #1

Let ,

X= Quantity Demand

P(x)= Probability

To find the No. of issues of the magazine that newsstand owner can expect to sell per week as,

E(x)=(10*0.20)+(11*0.15)+(12*0.30)+(13*0.10)+(14*0.05)+(15*0.20)

E(x)=2+1.65+3.6+1.3+0.7+3=12.25

Ans: The No. of issues of the magazine that newsstand owner can expect to sell per week is 12.25.